# Using analytic models to describe effective PDFs

### v1

@@ Abstract
Parton distribution functions (PDFs) play a pivotal role in hadron collider phenomenology. They are non-perturbative quantities extracted from fits to available data, and their scale dependence is dictated by the DGLAP evolution equations. In this article, we discuss machine learning (ML) strategies to efficiently compute PDFs by bypassing the scale evolution. Analytical approximations to the PDFs as a function of $x$ and $Q^{2}$, including up to next-to-leading order (NLO) QCD effects, are obtained. The methodology is tested by reproducing the HERAPDF2.0 set and implementing the analytical expressions in benchmarking codes. It is found that the computational cost is reduced while the precision of the simulations stays well under control.
@@ I Introduction
Most of the current higher-order cross-section simulations rely on the factorization theorem Collins et al. (1989); Catani et al. (2012); Cieri et al. (2024), i.e., splitting the whole calculation into perturbative and non-perturbative sectors. The process-dependent perturbative contribution is given in terms of well-defined analytical functions (or highly-efficient numerical representations of them). The universal, non-perturbative part, encoded in parton distribution functions (PDFs), can not be computed from fist principles: only their evolution with the factorization scale chosen can be calculated via the DGLAP Altarelli and Parisi (1977) evolution equations. The determination of the PDFs is, instead, achieved by performing global fits to existing data. For practical implementations, the outcome of these fits are given as tables in the relevant kinematic variables with a function that interpolates over them. The interpolation requires to perform several evaluations and the routines are written to be as fast as possible to allow for the efficient computation of physical observables. However, some cross-sections might require a very large number of trials (and running time) to achieve relevant precision, due to the nature of the process. Thus, it is worth exploring the possibility of reducing the CPU/GPU time required by providing an alternative form for the PDFs. Therefore, in this work, we aim to determine an analytical functional form for a set of known PDFs, by approximating their $(x,Q^{2})$-behaviour. This is equivalent to identify approximations to closed analytic solutions of DGLAP equations, a highly non-trivial problem in the context of coupled integro-differential equations Simonelli (2024).
@@
To achieve this ambitious objective, we profit from artificial intelligence. By using Machine-Learning (ML) techniques, we identify suitable functional dependencies to accurately describe the PDF sets. Our thesis is that, avoiding the interpolation over the grids, it will drastically reduce the time required to perform the simulations/computations. Furthermore, the PDFs would be then written in terms of a few parameters, reducing the storage required for them. Also, having access to analytic formulae for PDFs will allow to calculate their derivatives, both in $x$ and $Q$, which are relevant for performing the matching of $F_{2}$ in the threshold regions.
@@
Throughout this work we focus on the set of collinear proton PDFs extracted by the xFitter collaboration, HERAPDF2.0 Abramowicz et al. (2015), though the methodology is in principle applicable to any other PDF set, provided it is determined via a similar procedure. We stress that our study does not replace, in any way, the need of performing global fits. On the contrary, it is only by having the PDF sets already determined that we can search a posteriori for a functional form.
@@
This work is organised as follows. In Sec. II we present our proposed analytical model for a set of collinear, unpolarised proton PDFs, using HERAPDF2.0 as baseline. The methodology used to determine the functional form is explained in detail in Sec. III. Our results and their comparison with the baseline PDFs can be found in Sec. IV. We also check, in Sec. V, the time required to run benchmarking codes using both the available PDF grids and our a posteriori analytical PDFs. We conclude by summarising our findings in Sec. VI.
@@ II Starting hypothesis
Traditionally, proton PDFs  [^1] This also applies, with some differences, to other non-perturbative quantities feasible of phenomenological studies in pQCD, such as fragmentation functions, fracture functions, transverse momentum dependent distributions, etc. are determined by proposing a functional form in Bjorken $x$ for the parton densities (or a linear combination of them) at some chosen initial scale $Q_{0}^{2}$. Giving values to the parameters involved and evolving the PDFs to the experimental scales $Q^{2}$ using the DGLAP equations, one obtains the distributions that, convoluted with the partonic cross-sections, is compared with data. Repeating the procedure until an adequate description is achieved (according to some chosen criteria), the best fit parameters are determined. This is the case of many sets of proton and nuclear PDFs (e.g. Bailey et al. (2021); Hou et al. (2021); Alekhin et al. (2017); Helenius et al. (2022); de Florian et al. (2012); Eskola et al. (2022); Duwentäster et al. (2021)), and fragmentation functions (e.g.Borsa et al. (2022, 2024)). Other possibility that does not use a proposed parametrization but rather relies on Neural Networks is also employed by the NNPDF collaboration (see e.g. Ref. Ball et al. (2022) and references within).
@@
To determine an analytical expression for the collinear proton PDFs, which we test with the HERAPDF2.0 set at NLO QCD accuracy, we propose that the $Q^{2}$ dependence of the PDFs is given by an extension of Eq. (1), with each parameter acquiring a $Q^{2}$ dependence. In other words, we start from the assumption that
@@ III Methodology
In order to obtain the coefficients from Eq. (2), we performed a two-step fitting procedure. First, we fixed a PDF set within the LHAPDF framework: in our study, it was HERAPDF20__\__NLO__\__EIG, obtained by the xFitter group, which uses solely data from the HERA collider. Then, we generated a grid of $5000-10000$ random points in $\{x,Q\}$ with $x\in[10^{-4},0.65]$ and $Q\in[Q_{\rm Min},1000]$, evaluating the corresponding PDF. The range in $x$ was chosen in accordance to the recommendations done in Ref. Abramowicz et al. (2015). For the gluon and light quarks, we set $Q_{\rm Min}=Q_{0}\approx 1.37$ GeV, whilst we used $Q_{\rm Min}=1.5$ GeV and $Q_{\rm Min}=4.5$ GeV for the charm and bottom quark, respectively.
@@
We would like to highlight that this definition of the error is suitable to control the validity of the sum rules. Also, since the PDFs contribute to the hadronic cross-section through a convolution with the partonic cross-section, we observed that the integral error successfully provides an error estimation for the physical observables. We will return to this point in Sec. V.
@@ IV Results
In this section, we explicitly report the expressions for the analytic approximations to all the PDF flavours provided by HERAPDF2.0 at NLO. We fitted the $u_{v}$, $u$, $d_{v}$, $\bar{d}$, $g$, $s$, $c$ and $b$ distributions following the functional forms proposed in each of the following sub-sections. For each value of $Q$, we extracted the coefficients $\left\{A_{i},B_{i},\ldots\right\}$ performing a fit in $x$ with the Mathematica function NonlinearModelFit Research (2008). Then, we performed a second fit to determine the $Q$-dependence of these coefficients. In this step, we relied on ML techniques that provide suitable functional forms to build a basis. The functions obtained after this two-step fitting procedure, together with the full set of analytic ML-PDFs are publicly, available in an ancillary file Ochoa-Oregon et al. (2024). For this reason we refrain from presenting in the manuscript a table with the fitted parameters and functional forms.
IV.1 Up quark distributions
The optimal fit was found by considering $u_{v}$ and $u$ distributions. To this end, we proposed
$x\,u_{v}(x,Q^{2})$ $=$ $A_{u_{v}}(Q^{2})\,x^{B_{u_{v}}(Q^{2})}\,(1-x)^{C_{u_{v}}(Q^{2})}$
$\times$ $\big{[}1+D_{u_{v}}(Q^{2})x+E_{u_{v}}(Q^{2})x^{2}$
$+$ $F_{u_{v}}(Q^{2})x^{3}+G_{u_{v}}(Q^{2})x^{4}+H_{u_{v}}(Q^{2})x^{5}\big{]}\,,$
$x\,u(x,Q^{2})$ $=$ $A_{u}(Q^{2})\,x^{B_{u}(Q^{2})}\,(1-x)^{C_{u}(Q^{2})}\,,$
$\times$ $\big{[}1+D_{u}(Q^{2})x+E_{u}(Q^{2})x^{2}$
$+$ $F_{u}(Q^{2})x^{3}\big{]}\,,$
as functional forms to adjust, simultaneously, the dependence on $x$ and $Q^{2}$ for $u_{v}$ and $u$ distributions, respectively. The $\bar{u}$ distribution is then given by
$\bar{u}(x,Q^{2})=u(x,Q^{2})-u_{v}(x,Q^{2})\,.$ (14)
Refer to captionRefer to caption
In Fig. 1, we present a comparison between our analytic ML-PDF approximations to $x\,u_{v}(x,Q^{2})$ (upper plot) and $x\,\bar{u}(x,Q^{2})$ (lower plot) against HERAPDF20__\__NLO__\__EIG PDF set. We appreciate a very good agreement for different values of $Q$; in particular, we show results for $10$ GeV (blue), $100$ GeV (green) and $500$ GeV (orange).
Refer to captionRefer to caption
IV.2 Down quark distributions
$x\,d_{v}(x,Q^{2})$ $=$ $A_{d_{v}}(Q^{2})\,x^{B_{d_{v}}(Q^{2})}\,(1-x)^{C_{d_{v}}(Q^{2})}$
$\times$ $\left[1+D_{d_{v}}(Q^{2})x^{2}+E_{d_{v}}(Q^{2})x^{4}\right.$
$+$ $\left.F_{d_{v}}(Q^{2})x^{6}\right]\,,$
$x\,\bar{d}(x,Q^{2})$ $=$ $A_{\bar{d}}(Q^{2})\,x^{B_{\bar{d}}(Q^{2})}\,(1-x)^{C_{\bar{d}}(Q^{2})}$
$d(x,Q^{2})=d_{v}(x,Q^{2})+\bar{d}(x,Q^{2})\,.$ (17)
We show the comparison between our analytic ML-PDF approximations to $x\,d_{v}(x,Q^{2})$ (upper plot) and $x\,\bar{d}(x,Q^{2})$ (lower plot) w.r.t. HERAPDF20__\__NLO__\__EIG PDF set in Fig. 2. Again, the agreement is excellent for the complete range of $Q$ considered in this analysis.
IV.3 Gluon distribution
To obtain an approximation to the gluon distribution, which is more involved than the up and down quarks, we proposed a functional form inspired by Eq. (8). Concretely, we used
@@ IV.3 Gluon distribution
In Fig. 3, we show our analytic ML-PDF approximation to $x\,g(x,Q^{2})$ w.r.t. the corresponding gluon PDF from HERAPDF20__\__NLO__\__EIG. Even if small fluctuations occur at low-$x$, the overall agreement is very good. Furthermore, the agreement remains when choosing different values of $Q$, thanks to splitting the analysis into four regions.
@@ V Efficiency and quality benchmarks
In order to estimate the discrepancies between our analytic ML-PDFs and the original HERAPDF2.0 distributions in a phenomenologically-relevant way, we rely on our definition of integral error given in Eq. (10). In Fig. 7, we show that the integral error for almost all the distributions (i.e. $u_{v}$, $d_{v}$, $\bar{d}$, $c$, $b$ and $g$) is below $0.5\,\%$ in the $Q\in[Q_{0},1000\,{\rm GeV}]$ range. For the strange quark PDF, we appreciate a larger deviation at higher values of $Q$, reaching up to $1.5\,\%$ error for $Q\approx 1000$ GeV. Similarly, $\bar{u}$ has roughly $4\,\%$ error for $Q<700$ GeV, although it reaches up to $7\,\%$ for $Q\in[780\,{\rm GeV},1000\,{\rm GeV}]$. Despite these discrepancies, in what follows we will demonstrate that their impact in the calculation of physical observables is very limited.
Refer to captionRefer to caption
Figure 7: Integral error for our analytic ML-PDF approximations w.r.t. HERAPDF distributions. In the upper plot, we show up and down quark distributions, whilst $s$, $c$, $b$ and gluon PDF are presented in the lower plot. The integral error of $\bar{u}$ in the present plot is divided by a factor $10$ with the purpose of easing the visualization of the results.

### v2

@@ Abstract
Parton distribution functions play a pivotal role in hadron collider phenomenology. They are non-perturbative quantities extracted from fits to available data, and their scale dependence is dictated by the DGLAP evolution equations. In this article, we discuss machine-assisted strategies to efficiently compute PDFs directly incorporating the scale evolution without the need of separately solving DGLAP equations. Analytical approximations to the PDFs as a function of $x$ and $Q^{2}$, including up to next-to-leading order effects in Quantum Chromodynamics, are obtained. The methodology is tested by reproducing the HERAPDF2.0 set and implementing the analytical expressions in benchmarking codes. It is found that the computational cost is reduced while the precision of the simulations stays well under control.
@@ I Introduction
Most of the current higher-order cross-section simulations rely on the factorization theorem Collins et al. (1989); Catani et al. (2012); Cieri et al. (2024), i.e., splitting the whole calculation into perturbative and non-perturbative sectors. The process-dependent perturbative contribution is given in terms of well-defined analytical functions (or highly-efficient numerical representations of them). The universal, non-perturbative part, encoded in parton distribution functions (PDFs), can not be computed from first principles: only their evolution with the factorization scale chosen can be calculated via the DGLAP Altarelli and Parisi (1977) evolution equations. The determination of the PDFs is, instead, achieved by performing global fits to existing data. For practical implementations, the outcome of these fits are given as tables in the relevant kinematic variables with a function that interpolates over them. The interpolation requires to perform several evaluations and the routines are written to be as fast as possible to allow for the efficient computation of physical observables. However, some cross-sections might require a very large number of trials (and running time) to achieve relevant precision, due to the nature of the process. Thus, it is worth exploring the possibility of reducing the CPU/GPU time required by providing an alternative form for the PDFs. Therefore, in this work, we aim to determine an analytical functional form for a set of known PDFs, by approximating their $(x,Q^{2})$-behaviour. This is equivalent to identify approximations to closed analytic solutions of DGLAP equations, a highly non-trivial problem in the context of coupled integro-differential equations Simonelli (2024).
@@
To achieve this ambitious objective, we profit from artificial intelligence, machine-learning and similar tools implemented within Mathematica built-in functions. By using machine-assisted techniques, we identify suitable functional dependencies to accurately describe the PDF sets. Our thesis is that, avoiding the interpolation over the grids, it will drastically reduce the time required to perform the simulations/computations. Furthermore, the PDFs would be then written in terms of a few parameters, reducing the storage required for them. Also, having access to analytic formulae for PDFs will allow to calculate their derivatives, both in $x$ and $Q$, which are relevant for performing the matching of $F_{2}$ in the threshold regions.
@@
Throughout this work we focus on the set of collinear proton PDFs extracted by the xFitter collaboration, HERAPDF2.0 Abramowicz et al. (2015). We would like to highlight that our methodology is applicable to any PDF set, independently of the order at which it is calculated: if the PDF set exists, it is possible to find an analytic approximation with the techniques explained in this article. For this reason, we stress that our study does not replace, in any way, the need of performing global fits. On the contrary, it is only by having the PDF sets already determined that we can search a posteriori for a functional form. Our analytic approximations to a PDF set constitute a first proof-of-concept, and we show that it successfully passed several physically-motivated quality checks.
@@
This work is organised as follows. In Sec. II we present our proposed analytical model for a set of collinear, unpolarised proton PDFs, using HERAPDF2.0 as baseline. The methodology used to determine the functional form is explained in detail in Sec. III. Our results and their comparison with the baseline PDFs can be found in Sec. IV. In Sec. V, we carefully quantify the errors induced by our analytic approximations and check the validity of the sum rules. There, we also check the time required to run benchmarking codes using both the available PDF grids and our a posteriori analytical PDFs. We conclude by summarising our findings in Sec. VI.
@@ II Starting hypothesis
Traditionally, proton PDFs  [^1] This also applies, with some differences, to other non-perturbative quantities feasible of phenomenological studies in perturbative Quantum Chromodynamics (pQCD), such as fragmentation functions, fracture functions, transverse momentum dependent distributions, etc. are determined by proposing a functional form in Bjorken $x$ for the parton densities (or a linear combination of them) at some chosen initial scale $Q_{0}^{2}$. Giving values to the parameters involved and evolving the PDFs to the experimental scales $Q^{2}$ using the DGLAP equations, one obtains the distributions that, convoluted with the partonic cross-sections, is compared with data. Repeating the procedure until an adequate description is achieved (according to some chosen criteria), the best fit parameters are determined. This is the case of many sets of proton and nuclear PDFs (e.g. Bailey et al. (2021); Hou et al. (2021); Alekhin et al. (2017); Helenius et al. (2022); de Florian et al. (2012); Eskola et al. (2022); Duwentäster et al. (2021)), and fragmentation functions (e.g.Borsa et al. (2022, 2024)). Other possibility that does not use a proposed parametrization but rather relies on Neural Networks is also employed by the NNPDF collaboration (see e.g. Ref. Ball et al. (2022) and references within).
@@
To determine an analytical expression for the collinear proton PDFs, which we test with the HERAPDF2.0 set at next-to-leading order (NLO) accuracy in QCD, we propose that the $Q^{2}$ dependence of the PDFs is given by an extension of Eq. (1), with each parameter acquiring a $Q^{2}$ dependence. In other words, we start from the assumption that
@@ III Methodology
In order to obtain the coefficients from Eq. (2), we performed a two-step fitting procedure. First, we fixed a PDF set within the LHAPDF framework: in our study, it was HERAPDF20__\__NLO__\__EIG, obtained by the xFitter group, which uses solely data from the HERA collider. Then, we generated a grid of ${\cal O}(5000)$ random points in $\{x,Q\}$ with $x\in[10^{-4},0.65]$ and $Q\in[Q_{\rm Min},1000]$, evaluating the corresponding PDF in these points. The range in $x$ was chosen in accordance to the recommendations done in Ref. Abramowicz et al. (2015). For the gluon and light quarks, we set $Q_{\rm Min}=Q_{0}\approx 1.37$ GeV, whilst we used $Q_{\rm Min}=1.5$ GeV and $Q_{\rm Min}=4.5$ GeV for the charm and bottom quark, respectively.
@@
We would like to highlight that this definition of the error is suitable to control the validity of the sum rules. Also, since the PDFs contribute to the hadronic cross-section through a convolution with the partonic cross-section, we observed that the integral error successfully provides an error estimation for the physical observables. Both tests are discussed extensively in Sec. V.
Besides this error definition, we explored additional estimators, such as the error shape:
$\tilde{\Delta}_{i}(Q^{2})=\frac{1}{N}\,\sum_{j=1}^{N}\Bigg{|}1-\frac{f^{\rm ML}_{i}(x_{j},Q^{2})}{f^{\rm HERA}_{i}(x_{j},Q^{2})}\Bigg{|}\,,$ (12)
where $\{x_{j}\}_{j=1,\ldots,N}$ is a partition of $x\in[10^{-4},1]$. This definition is particularly sensitive to fluctuations or oscillations around the original PDF, and leads to a noticeable overestimation of the error. For example, we found typical error shapes of ${\cal O}(10-50\,\%)$ for a large range of $Q$ values, in spite of percent and even sub-percent level deviations of the sum rules and other physical observables (see Sec. V). For this reason, we consider the integral error as a more reliable estimator.
@@ IV Results
In this section, we explicitly report the expressions for the analytic approximations to all the PDF flavours provided by HERAPDF2.0 at NLO. We fitted the $u_{v}$, $\bar{u}$, $d_{v}$, $\bar{d}$, $g$, $s$, $c$ and $b$ distributions following the functional forms proposed in each of the following sub-sections. For each value of $Q$, we extracted the coefficients $\left\{A_{i},B_{i},\ldots\right\}$ performing a fit in $x$ with the Mathematica built-in function NonlinearModelFit Research (2008).
Then, we performed a second fit to determine the $Q$-dependence of these coefficients. In this step, we relied on machine-assisted techniques that provide suitable functional forms to build the ansatzes. In particular, we make use of the Mathematica built-in routine FindFormula to generate an approximation to the coefficients $\left\{A_{i},B_{i},\ldots\right\}$ as functions of $Q$. Then, we used these approximations to propose an ansatz for each coefficient, including similar expressions to those found by FindFormula. The functions obtained after this two-step fitting procedure, together with the full set of analytic ML-PDFs are publicly available in an ancillary file uploaded to the Zenodo repository Ochoa-Oregon et al. (2024). Given that the final number of parameters is quite large, we refrain from presenting in the manuscript a table with the fitted parameters and functional forms.
IV.1 Gluon distribution
To obtain an approximation to the gluon distribution, we proposed a functional form inspired by Eq. (8). Concretely, we used
@@ IV.1 Gluon distribution
In Fig. 1, we show our analytic ML-PDF approximation to $x\,g(x,Q^{2})$ w.r.t. the corresponding gluon PDF from HERAPDF20__\__NLO__\__EIG, at $10$ GeV (blue), $100$ GeV (green) and $1000$ GeV (orange). Even if small fluctuations occur at low-$x$, the overall agreement is very good. Furthermore, the agreement remains when choosing different values of $Q$, thanks to splitting the analysis into four regions.
IV.2 Down quark distributions
$x\,d_{v}(x,Q^{2})$ $=$ $A_{d_{v}}(Q^{2})\,x^{B_{d_{v}}(Q^{2})}\,(1-x)^{C_{d_{v}}(Q^{2})}$
$\times$ $\left[1+D_{d_{v}}(Q^{2})x^{2}+E_{d_{v}}(Q^{2})x^{4}\right.$
$+$ $\left.F_{d_{v}}(Q^{2})x^{6}\right]\,,$
$x\,\bar{d}(x,Q^{2})$ $=$ $A_{\bar{d}}(Q^{2})\,x^{B_{\bar{d}}(Q^{2})}\,(1-x)^{C_{\bar{d}}(Q^{2})}$
$d(x,Q^{2})=d_{v}(x,Q^{2})+\bar{d}(x,Q^{2})\,.$ (22)
In Fig. 2, we present a comparison between our analytic ML-PDF approximations to $x\,d_{v}(x,Q^{2})$ (upper plot) and $x\,\bar{d}(x,Q^{2})$ (lower plot) and the HERAPDF20__\__NLO__\__EIG PDF set. We appreciate a very good agreement for different values of $Q$.
IV.3 Up quark distributions
The optimal fit was found by considering $u_{v}$ and $\bar{u}$ distributions. To this end, we proposed
$x\,u_{v}(x,Q^{2})$ $=$ $A_{u_{v}}(Q^{2})\,x^{B_{u_{v}}(Q^{2})}\,(1-x)^{C_{u_{v}}(Q^{2})}$
$\times$ $\big{[}1+D_{u_{v}}(Q^{2})x+E_{u_{v}}(Q^{2})x^{2}$
$+$ $F_{u_{v}}(Q^{2})x^{3}+G_{u_{v}}(Q^{2})x^{4}+H_{u_{v}}(Q^{2})x^{5}\big{]}\,,$
to approximate $u_{v}$, whilst a slightly more complicated ansatz was used for $\bar{u}$. In that case, we split the analysis into two regions:
$R_{1}$ $=$ $\{Q_{0}\leq Q\leq 10\,{\rm GeV}\}\,,$ (24)
$R_{2}$ $=$ $\{10\,{\rm GeV}\leq Q\leq 1000\,{\rm GeV}\}\,.$ (25)
Then, we defined
$x\,\bar{u}(x,Q^{2})$ $=$ $A_{\bar{u}}(Q^{2})\,x^{B_{\bar{u}}(Q^{2})}\,(1-x)^{C_{\bar{u}}(Q^{2})}$
$\times$ $\big{[}1+D_{\bar{u}}(Q^{2})x+E_{\bar{u}}(Q^{2})x^{2}$
$+$ $F_{\bar{u}}(Q^{2})x^{3}+G_{\bar{u}}(Q^{2})x^{4}+H_{\bar{u}}(Q^{2})x^{5}\big{]}\,,$
for $Q\in R_{1}$ and
$x\,\bar{u}(x,Q^{2})$ $=$ $A_{\bar{u}}(Q^{2})\,x^{B_{\bar{u}}(Q^{2})}\,(1-x)^{C_{\bar{u}}(Q^{2})}$ (26)
$\times$ $\big{[}1+D_{\bar{u}}(Q^{2})x+E_{\bar{u}}(Q^{2})x^{3}\big{]}$
-- $\Theta(x_{C,\bar{u}}-x)A^{\prime}_{\bar{u}}(Q^{2})\,x^{B^{\prime}_{\bar{u}}(Q^{2})}$
$\times$ $(1-x)^{C^{\prime}_{\bar{u}}(Q^{2})}[1+D^{\prime}_{\bar{u}}(Q^{2})x^{2}]\,,$
for $Q\in R_{2}$. Here, we used the cut $x_{C,\bar{d}}=0.01$, inspired by the gluon PDF parametrization defined by HERAPDF. After fitting these two distributions, we can define
$u(x,Q^{2})=\bar{u}(x,Q^{2})+u_{v}(x,Q^{2})\,,$ (27)
and obtain the $u$-quark distribution.
We show the comparison between our analytic ML-PDF approximations to $x\,u_{v}(x,Q^{2})$ (upper plot) and $x\,\bar{u}(x,Q^{2})$ (lower plot) w.r.t. HERAPDF20__\__NLO__\__EIG PDF set in Fig. 3. Again, the agreement is excellent for the complete range of $Q$ considered in this analysis.
@@ V Efficiency and quality benchmarks
In order to estimate the discrepancies between our analytic ML-PDFs and the original HERAPDF2.0 distributions in a phenomenologically-relevant way, we rely on our definition of integral error given in Eq. (10). In Fig. 7, we show that the integral error for almost all the distributions (i.e. $u_{v}$, $d_{v}$, $\bar{u}$, $\bar{d}$, $c$, $b$ and $g$) is below $0.5\,\%$ in the $Q\in[Q_{0},1000\,{\rm GeV}]$ range. For the strange quark PDF, we appreciate a larger deviation at higher values of $Q$, reaching up to $1.5\,\%$ error for $Q\approx 1000$ GeV. Although not present in the plots, we want to highlight that the integral error for $u$ and $d$ is also well under control, being below percent level for the whole range of $Q$-values explored. This is expected from Eqs. (22) and (27), propagating the errors shown in Fig. 7.
Figure 7: Integral error for our analytic ML-PDF approximations w.r.t. HERAPDF distributions. In the upper plot, we show up and down quark distributions, whilst $s$, $c$, $b$ and gluon PDF are presented in the lower plot.
As we explained in Sec. III, the integral error is expected to provide a more reliable estimation on the impact of using our ML-PDFs instead of the original PDF set in a physical calculation. Still, for the sake of completeness, we present here an analysis of the approximation error as a function of $x$. For this purpose, we sample with an exponential distribution the range $Q\in[10\,{\rm GeV},1000\,{\rm GeV}]$ using $N=50,000$ random points. Then, we define
$\bar{\Delta}_{i}(x)=\frac{1}{N}\sum_{j=1}^{N}\,\Bigg{|}1-\frac{f_{i}^{\rm ML}(x,Q_{j}^{2})}{f_{i}^{\rm HERA}(x,Q_{j}^{2})}\Bigg{|}\,,$ (40)
as an estimator of the error in $x$. This definition is similar to the error shape mentioned in Sec. III, and corresponds to an average of the relative errors as a function of $x$. In Fig. 8, we show the results for up, down (upper plot), strange, charm, bottom quarks and the gluon (lower plot). For $u_{v}$, $d_{v}$, $\bar{u}$ and $\bar{d}$-quarks, $\bar{\Delta}$ is below $10\,\%$ for $x\in[10^{-3},0.3]$. Similarly, for the other quark flavours, the error is below $20\,\%$ in $x\in[10^{-3},0.3]$. For the gluon, the error is slightly larger, ${\cal O}(10-25\,\%)$ in the central $x$ region. All the distributions tend to increase their average relative error for $x>0.3-0.4$, since the PDFs decrease a couple of orders of magnitude for larger values of $x$. Also, the valence distributions show this behavior in the low-$x$ region. In both cases, small absolute discrepancies translate into large relative fluctuations. This effect is shadowed in the definition of integral error, because the contribution of the PDFs in those regions are rather small.
Another important observation is the presence of oscillations in the relative error in the region $x\approx 0.05-0.1$, particularly for distributions that made use of a gluon-like functional form. This behaviour is due to the matching of two Euler beta functions around $x_{C}={\cal O}(0.1)$, which originates a fluctuation since they have different signs. In fact, this justifies the reduced integral error because these fluctuations around the original PDF cancel at integrand level.
Figure 8: Estimation of the error as a function of $x$, averaged over $Q\in[10\,{\rm GeV},1000\,{\rm GeV}]$ for our analytic ML-PDF approximations w.r.t. HERAPDF2.0 distributions. In the upper plot, we show up and down quark distributions, whilst $s$, $c$, $b$ and gluon PDFs are presented in the lower plot.